OpenStudy (anonymous):

2a + b = 5 a - b = 4 what is the answer????

OpenStudy (anonymous):

a = 3, b = -1

OpenStudy (hoblos):

3a=9 a=3 2a + b = 5-----> b=5-6 b=-1